OpenStudy (anonymous):

Graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (1,-2),m=-3/5 Can anyone help??? I need to see the step by step instructions because I do not understand how to do these type problems. Would anyone be willing to share step by step instructions?

OpenStudy (anonymous):

y + 2 = -3/5 (x-1) now find the equation

OpenStudy (anonymous):

form: y - y_1 = m( x - x_1)